# Towards Better Code Generation: Adaptive Decoding with Uncertainty Guidance

### v2

@@ III Approach
III-A Overview
Figure 4: Approach Overview of AdaDec
Given a target model LM, the adaptive decoding process of AdaDec, which incorporates the pause-and-rerank mechanism, is illustrated in Figure 4. At each decoding step $t$, after LM predicts a probability distribution $\bm{p}_{t}$ for the next token, AdaDec determines whether to pause the decoding by checking if the entropy $H(\bm{p}_{t})$ of $\bm{p}_{t}$ exceeds a model-specific threshold $\tau^{\bm{LM}}$. If $H(\bm{p}_{t})$ surpasses $\tau^{\bm{LM}}$, indicating high uncertainty, AdaDec triggers the reranking process. Specifically, it selects the top-$K$ candidate tokens from the distribution and applies a lookahead strategy, re-scoring each token’s probability based on its expected likelihood over the next $N$ decoding steps. The token with the highest re-scored probability is then selected as the next token. This process continues iteratively until an end-of-sequence (EOS) token is generated.
@@ III-A Overview
III-B Entropy-based Pause Checking
@@ III-B1 Data Collection
We generate training data by conducting ground-truth–guided next-token prediction with the target model LM on the BigCodeBench dataset [22]. In this process, the ground truth token is provided as input at each step instead of the model’s own prediction. During this process, we record the model’s predicted probability distribution over the vocabulary at each decoding step and calculate the entropy of the predicted distribution. Based on whether the ground truth token is ranked as the top-1 prediction, we assign a binary label: 1 if it is top-1, and 0 otherwise. Based on this process, we collect between approximately 100,000 and 160,000 positive samples (labeled as 1) and between 17,000 and 23,000 negative samples (labeled as 0) per model. For example, in the case of DS-6.7B, we obtain 161,856 positive samples and 19,023 negative samples. The substantial imbalance between the number of positive and negative samples further highlights the inefficiency of beam search, which performs reranking at every decoding step regardless of model uncertainty.
We then balance the dataset by randomly downsampling the positive samples and selecting an equal number of negative samples, which is a common technique for handling imbalanced classification tasks [23]. This choice is driven by the nature of logic drift in code generation: once a drift point occurs, it typically leads to irreversible semantic errors. Therefore, it is essential to maximize the classifier’s ability to detect negative cases, even if they are relatively scarce. By constructing a balanced dataset, we prevent the model from being biased toward the majority class and enable it to more effectively learn the decision boundary for identifying uncertain decoding steps. This strategy enhances the classifier’s sensitivity to potential drift points, thereby improving the overall reliability of the entropy-based triggering mechanism.
III-B2 Regression Model Training
@@ III-B4 Pause Triggering
III-C Lookahead-based Token Reranking
AdaDec employs a lookahead re-scoring strategy for token reranking, enhancing selection by assessing the future impact of each candidate token at every decoding step. Inspired by A* search heuristics and prior work on constrained generation [17], this method extends token probability evaluation beyond immediate next-token predictions to consider longer-term sequence quality. The process is governed by two hyperparameters: the lookahead beam size ($B$), which determines the number of candidate tokens considered, and the lookahead length ($L$), which sets the length of the future token sequence evaluated.
III-C1 Lookahead Process
@@
where $P(\cdot)$ denotes the model’s predicted probability for a token. To ensure a fair comparison across candidates with varying lookahead lengths (as discussed in the next section on lookahead length selection), we normalize the total log-probability by computing the geometric mean of token probabilities in the above equation, rather than relying solely on log-likelihood. This normalization mitigates the bias against longer sequences, which tend to accumulate lower joint probabilities. The resulting score captures the model’s overall confidence in the full trajectory, balancing immediate token likelihood with the coherence of future predictions.
@@
III-C2 Lookahead Beam Size Selection
To determine an appropriate value for $B$, we analyze the training data collected in Section III-B1. Specifically, we examine the average rank of the ground truth token at decoding steps where the Shannon entropy is either above or below the threshold $\tau^{\bm{LM}}$, which is learned via logistic regression. After filtering out samples where the ground truth token’s rank exceeds 20, we observe a clear distinction: when the entropy is below or equal to the threshold, the average rank typically falls between 1.1 and 1.3, indicating high model confidence. In contrast, when the entropy exceeds the threshold—thus triggering a pause—the average rank increases to between 2.6 and 3.4.
@@ III-C2 Lookahead Beam Size Selection
III-C3 Lookahead Length Selection
We explore two strategies for determining the lookahead length $L$ during trajectory scoring:
  • Fixed: A simple approach that uses a constant lookahead length (e.g., 5) for all candidate tokens.
[…]
The line-aware dynamic strategy allows for adaptive lookahead tailored to the syntactic structure of code, potentially improving alignment with natural coding boundaries. In Section IV, we empirically compare the two strategies.
IV Evaluation
@@
  • RQ3 (Ablation Study): How effective is the logistic regression–based threshold learning component, and how do the two lookahead length selection strategies affect performance?
  • RQ4 (Generalizability): How well does AdaDec generalize across different models and datasets?
IV-A Experimental Setup
@@
Datasets. We evaluate our method on two widely used code generation benchmarks. The first is HumanEval [16], which contains 164 hand-crafted problems. The second is MBPP [26], from which we use the first 200 problems for evaluation.
Baselines. We compare AdaDec against two standard decoding strategies: Greedy decoding and Beam Search with a beam size of 3 (equal to the lookahead beam size used in AdaDec). These baselines represent common trade-offs between generation quality and computational efficiency.
Variants of AdaDec. To examine the impact of different lookahead length (L) strategies, we evaluate two variants of our method:
  • $\textsc{AdaDec}_{L=5}$: Applies the fixed lookahead length of $L=5$ for all reranking steps.
  • $\textsc{AdaDec}_{dynL}$: Uses the dynamic, line-aware lookahead length with a limit of 10.
IV-B RQ1: Effectiveness in Improving Functional Correctness
@@
Table II presents the Pass@1 performance of four decoding strategies, Greedy, Beam Search, $\textsc{AdaDec}_{L=5}$, and $\textsc{AdaDec}_{dynL}$, across eight LLMs evaluated on the HumanEval and MBPP datasets.
@@
Table II: Pass@1 comparison of Greedy, Beam Search, $\textsc{AdaDec}_{L=5}$, and $\textsc{AdaDec}_{dynL}$ on HumanEval and MBPP. Values in parentheses indicate the change relative to Greedy. The “AVG. $\Delta$” row reports the average difference from Greedy across all eight models.
Model HumanEval MBPP
Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$ Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$
DS-1.3B 56.71% 59.76% (+3.05%) 64.02% (+7.32%) 61.59% (+4.88%) 36.50% 40.50% (+4.00%) 37.00% (+0.50%) 37.00% (+0.50%)
DS-6.7B 71.95% 74.39% (+2.44%) 74.39% (+2.44%) 70.73% (-1.22%) 47.50% 51.00% (+3.50%) 51.00% (+3.50%) 52.50% (+5.00%)
ST-3B 55.49% 55.49% (+0.00%) 55.49% (+0.00%) 55.49% (+0.00%) 35.00% 37.50% (+2.50%) 40.00% (+5.00%) 40.00% (+5.00%)
CL-7B 38.41% 42.68% (+4.27%) 40.24% (+1.83%) 42.68% (+4.27%) 35.00% 36.50% (+1.50%) 38.50% (+3.50%) 40.00% (+5.00%)
QW-0.6B 21.34% 28.05% (+6.71%) 25.00% (+3.66%) 25.61% (+4.27%) 12.00% 20.50% (+8.50%) 18.00% (+6.00%) 19.00% (+7.00%)
QW-1.7B 46.34% 46.34% (+0.00%) 50.00% (+3.66%) 47.56% (+1.22%) 34.00% 40.00% (+6.00%) 37.00% (+3.00%) 37.00% (+3.00%)
QW-4B 57.32% 65.85% (+8.54%) 64.02% (+6.71%) 67.07% (+9.76%) 23.50% 35.00% (+11.50%) 39.00% (+15.50%) 37.00% (+13.50%)
QW-8B 61.59% 66.46% (+4.88%) 71.34% (+9.76%) 70.12% (+8.54%) 40.50% 49.00% (+8.50%) 49.00% (+8.50%) 50.50% (+10.00%)
AVG. $\Delta$ +3.73% +4.42% +3.96% +5.75% +5.69% +6.12%
Comparison with Greedy Search. Both variants of our adaptive decoding approach, $\textsc{AdaDec}_{L=5}$ and $\textsc{AdaDec}_{dynL}$, consistently outperform Greedy decoding across nearly all models and datasets, with the exception of $\textsc{AdaDec}_{dynL}$ on DS-6.7B for HumanEval. On HumanEval, $\textsc{AdaDec}_{L=5}$ yields an average absolute improvement of +4.42%, while $\textsc{AdaDec}_{dynL}$ follows closely with +3.96%, compared to the Greedy baseline. The performance gains are even more prominent on MBPP, where $\textsc{AdaDec}_{dynL}$ achieves the highest average improvement of +6.12%, with $\textsc{AdaDec}_{L=5}$ closely behind at +5.69%.
Notably, certain models show substantial improvements. For instance, QW-8B exhibits a +9.76% increase with $\textsc{AdaDec}_{L=5}$ on HumanEval and a +10.00% increase with $\textsc{AdaDec}_{dynL}$ on MBPP. Similarly, QW-4B sees large gains across both datasets (ranging from +6.71% to +15.50%), underscoring the effectiveness of adaptive decoding especially for mid- to large-sized models.
Comparison with Beam Search. When compared to Beam Search, the performance of our adaptive decoding strategies is competitive and often superior, particularly on MBPP. On HumanEval, $\textsc{AdaDec}_{L=5}$ slightly outperforms Beam Search on average (+4.42% vs. +3.73% over Greedy), while $\textsc{AdaDec}_{dynL}$ achieves comparable gains (+3.96%). On MBPP, $\textsc{AdaDec}_{dynL}$ demonstrates the best average improvement over Greedy (+6.12%), surpassing both Beam Search (+5.75%) and $\textsc{AdaDec}_{L=5}$ (+5.69%).
While Beam Search excels in some individual cases (e.g., DS-6.7B and CL-7B on HumanEval), adaptive decoding methods show more robust performance across both datasets and models. $\textsc{AdaDec}_{L=5}$ and $\textsc{AdaDec}_{dynL}$ each achieve the highest Pass@1 scores in 6 out of the 16 evaluation settings (2 datasets × 8 models), highlighting their consistent generalization advantage across tasks and model sizes. Moreover, our adaptive decoding approach offers improved efficiency, as discussed in the next section (RQ2).
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Summary: Our uncertainty-guided adaptive decoding approach, AdaDec, outperforms Greedy decoding on average by up to +6.12% in Pass@1 across HumanEval and MBPP, and achieves competitive or superior results compared to Beam Search. $\textsc{AdaDec}_{dynL}$ delivers the highest gains on MBPP, while $\textsc{AdaDec}_{L=5}$ leads on HumanEval.
IV-C RQ2: Efficiency in Computational Cost
@@
We evaluate the efficiency of AdaDec in terms of computational cost on the HumanEval and MBPP datasets from the following perspectives. First, we measure the pause rate of AdaDec to assess the possible overhead introduced by its pause-then-rank mechanism. Second, we compare the count of generated tokens and the prediction latency of AdaDec against Greedy decoding and Beam Search.
@@
Pause Rate. As presented in Table III, the pause rate ranges from 3.95% to 11.88%, with an average of 7.15%. This low activation frequency indicates that AdaDec concentrates computational effort on a small subset of high-uncertainty steps, effectively balancing local correction with overall efficiency. In all experiments, the pause rate remains below 12%, demonstrating that selective triggering successfully avoids the high cost associated with global lookahead.
Table III: Pause ratio of the models on HumanEval and MBPP.
DS-1.3B DS-6.7B ST-3B CL-7B QW-0.6B QW-1.7B QW-4B QW-8B
$\textsc{AdaDec}_{L=5}$
HumanEval 6.54% 5.53% 4.76% 3.95% 6.10% 11.07% 7.95% 5.98%
MBPP 8.32% 5.82% 6.35% 5.20% 6.33% 11.41% 10.64% 8.41%
$\textsc{AdaDec}_{dynL}$
HumanEval 6.69% 5.60% 4.72% 3.95% 5.18% 10.06% 9.37% 6.39%
MBPP 8.28% 5.89% 6.51% 5.45% 6.21% 11.88% 10.18% 7.96%
Token Count. Table IV summarizes the results. A consistent and noteworthy outcome across both datasets is that AdaDec substantially reduces the average number of generated tokens. This reduction is not achieved by truncating outputs or constraining sequence length, but rather through more focused and efficient content selection during decoding. By leveraging its pause-and-rank mechanism to identify promising continuations early, AdaDec avoids veering into verbose sequences.
Compared with Greedy decoding, representative examples include: on MBPP, QW-0.6B reduces token count by 54.72%, from 341.07 to 154.45 with $\textsc{AdaDec}_{dynL}$; on HumanEval, CL-7B drops from 104.92 to 88.43 tokens (-15.71%) with $\textsc{AdaDec}_{dynL}$; and on MBPP, both ST-3B and QW-1.7B achieve 24–34% reduction swith $\textsc{AdaDec}_{L=5}$. These results highlight AdaDec’s built-in “generation compression” effect: by pausing to rank and select more meaningful continuations, the model often generates correct code using fewer tokens.
Latency. As shown in Table V, AdaDec sometimes reduces total generation latency, despite the additional overhead of pauses and candidate reranking. This counterintuitive outcome stems from the shorter outputs produced, which in many cases offset, or even outweigh, the cost of reranking. On MBPP, for instance, $\textsc{AdaDec}_{L=5}$ reduces the average latency for Qwen3-0.6B by 31.12% (from 6.92 s to 4.76 s), for Stable-3B by 15.53% (1.17 s to 0.99 s), and for CodeLlama-7B by 18.10% (3.23 s to 2.65 s). On HumanEval, $\textsc{AdaDec}_{dynL}$ achieves a 13.45% reduction for Qwen3-0.6B (3.67 s to 3.17 s).
Regarding Beam Search, although its latency is comparable to AdaDec, its computational cost is substantially higher. This is because Beam Search performs many more logits computations and token predictions during decoding. While these computations are parallelized on CPUs or GPUs and thus do not significantly increase latency, the added computational overhead results in greater energy consumption.
Table IV: Comparison of the average number of tokens generated per problem by Greedy, Beam Search, $\textsc{AdaDec}_{L=5}$, and $\textsc{AdaDec}_{dynL}$ on HumanEval and MBPP (the brackets indicate the change ratio relative to Greedy)
Model HumanEval MBPP
Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$ Greedy Beam Search $\textsc{AdaDec}_{L=5}$ $\textsc{AdaDec}_{dynL}$
DS-1.3B 66.27 74.80 (+12.86%) 64.38 (-2.86%) 66.59 (+0.47%) 75.61 71.10 (-5.96%) 54.49 (-27.93%) 59.36 (-21.49%)
DS-6.7B 64.32 67.68 (+5.21%) 62.10 (-3.45%) 60.76 (-5.55%) 70.25 62.18 (-11.48%) 59.28 (-15.62%) 53.54 (-23.78%)
ST-3B 79.98 73.23 (-8.44%) 70.18 (-12.25%) 68.10 (-14.85%) 71.83 63.12 (-12.13%) 54.66 (-23.91%) 53.53 (-25.48%)
CL-7B 104.92 104.65 (-0.26%) 90.04 (-14.19%) 88.43 (-15.71%) 90.88 81.03 (-10.83%) 68.21 (-24.94%) 68.61 (-24.50%)
QW-0.6B 168.62 98.71 (-41.46%) 107.54 (-36.22%) 100.10 (-40.64%) 341.07 80.94 (-76.27%) 171.21 (-49.80%) 154.45 (-54.72%)
QW-1.7B 100.57 73.57 (-26.84%) 81.89 (-18.57%) 83.40 (-17.07%) 131.80 85.45 (-35.17%) 86.38 (-34.46%) 98.66 (-25.15%)
QW-4B 115.34 89.61 (-22.31%) 111.87 (-3.01%) 89.08 (-22.77%) 86.55 52.50 (-39.35%) 82.72 (-4.43%) 74.22 (-14.25%)
QW-8B 95.24 81.85 (-14.06%) 84.79 (-10.97%) 89.10 (-6.45%) 94.58 60.25 (-36.30%) 78.34 (-17.18%) 71.96 (-23.91%)

### v3

@@ 3. Approach
3.1. Overview
Figure 5. Approach Overview of AdaDec
As illustrated in Figure 5, AdaDec is an adaptive (dynamic) decoding framework: at inference time it continuously monitors per-step uncertainty and dynamically decides whether to intervene, rather than applying a uniform policy across all steps or only at line-start positions. Given a target model LM, at each decoding step $t$ the model produces a next-token distribution $\bm{p}_{t}$. AdaDec computes the Shannon entropy $H(\bm{p}_{t})$ and compares it against a model-specific threshold $\tau^{\bm{LM}}$. If $H(\bm{p}_{t})\leq\tau^{\bm{LM}}$, the framework proceeds with the base decoding policy (e.g., greedy, top-$k$, or top-$p$ sampling). If $H(\bm{p}_{t})>\tau^{\bm{LM}}$, AdaDec pauses normal decoding and triggers a targeted reranking routine over the top-$K$ candidates. During reranking, each candidate undergoes a short lookahead of up to $N$ steps to estimate the quality of its future trajectory, and the candidate with the highest score is selected as the next token. This per-step decision loop repeats until an end-of-sequence (EOS) token is generated.
The adaptive nature of AdaDec has three practical consequences. First, the intervention is selective and data-driven: only high-uncertainty steps incur additional computation, which limits overhead compared to always-on strategies. Second, the threshold $\tau^{\bm{LM}}$ is learned in a preparation stage (we describe the learning procedure in Sec. 3.2) so that the pause decision is model-aware rather than hand-tuned. Third, the framework supports flexible reranking configurations (e.g., adjustable lookahead length $N$ and lookahead width $K$), allowing trade-offs between accuracy and latency to be tuned for different deployment constraints.
AdaDec is designed to be broadly applicable. Since the pause-then-rerank mechanism is triggered purely by uncertainty signals (entropy), it can be seamlessly integrated with various decoding strategies, including sampling. This flexibility underscores the generality of AdaDec: rather than being tied to a specific decoding style, it acts as a plug-in framework that can enhance both deterministic (e.g., greedy) and stochastic (e.g., sampling) generation. Unless otherwise specified, we use greedy decoding as the base policy throughout the remainder of this paper.
@@ 3.1. Overview
3.2. Entropy‐based Pause Checking
@@ 3.2.1. Data Collection
We generate training data by conducting ground-truth–guided next-token prediction with the target model LM on the BigCodeBench dataset (Zhuo et al., 2024), while carefully ensuring that the construction of our training data does not overlap with the evaluation benchmarks (HumanEval+, MBPP+, and DevEval+). To mitigate the risk of data contamination, we manually reviewed problem statements and function signatures to check for potential duplicates or near-duplicates between BigCodeBench and the evaluation sets. In this process, the ground truth token is provided as input at each step instead of the model’s own prediction. During this process, we record the model’s predicted probability distribution over the vocabulary at each decoding step and calculate the entropy of the predicted distribution. Based on whether the ground truth token is ranked as the top-1 prediction, we assign a binary label: 1 if it is top-1, and 0 otherwise. Based on this process, we collect between approximately 100,000 and 160,000 positive samples (labeled as 1) and between 17,000 and 23,000 negative samples (labeled as 0) per model. For example, in the case of DS-6.7B, we obtain 161,856 positive samples and 19,023 negative samples.
We then balance the dataset by randomly downsampling the positive samples and selecting an equal number of negative samples, which is a common technique for handling imbalanced classification tasks (Lemaitre et al., 2016). This choice is driven by the nature of logic drift in code generation: once a drift point occurs, it typically leads to irreversible semantic errors. Therefore, it is essential to maximize the classifier’s ability to detect negative cases, even if they are relatively scarce. By constructing a balanced dataset, we prevent the model from being biased toward the majority class and enable it to more effectively learn the decision boundary for identifying uncertain decoding steps. This strategy enhances the classifier’s sensitivity to potential drift points, thereby improving the overall reliability of the entropy-based triggering mechanism.
3.2.2. Regression Model Training
@@ 3.2.4. Pause Triggering
3.3. Lookahead-based Token Reranking
AdaDec employs a lookahead re-scoring strategy for token reranking, enhancing selection by assessing the future impact of each candidate token at every decoding step. Inspired by A* search heuristics and prior work on constrained generation (Lu et al., 2021), this method extends token probability evaluation beyond immediate next-token predictions to consider longer-term sequence quality. The process is governed by two hyperparameters: the lookahead beam size ($B$), which determines the number of candidate tokens considered, and the lookahead length ($L$), which sets the length of the future token sequence evaluated.
3.3.1. Lookahead Process
@@
where $P(\cdot)$ denotes the model’s predicted probability for a token. Because we apply an early stopping mechanism, different candidates may result in varying lookahead lengths. To ensure a fair comparison across such candidates, we normalize the total log-probability by computing the geometric mean of token probabilities in the above equation, rather than relying solely on log-likelihood. This normalization mitigates the bias against longer sequences, which tend to accumulate lower joint probabilities. The resulting score captures the model’s overall confidence in the full trajectory, balancing immediate token likelihood with the coherence of future predictions.
@@
3.3.2. Lookahead Beam Size Selection
To determine an appropriate value for $B$, we analyze the training data collected in Section 3.2.1. Specifically, we examine the average rank of the ground truth token at decoding steps where the Shannon entropy is either above or below the threshold $\tau^{\bm{LM}}$, which is learned via logistic regression. After filtering out samples where the ground truth token’s rank exceeds 20, we observe a clear distinction: when the entropy is below or equal to the threshold, the average rank typically falls between 1.1 and 1.3, indicating high model confidence. In contrast, when the entropy exceeds the threshold—thus triggering a pause—the average rank increases to between 2.6 and 3.4.
@@ 3.3.2. Lookahead Beam Size Selection
3.3.3. Lookahead Length Selection
We adopt a lookahead length parameter $L$ when performing trajectory scoring. At each reranking step we simulate up to $L$ future tokens for each candidate to estimate the quality of its continuation. The parameter $L$ controls the trade-off between accuracy and efficiency: a larger $L$ provides a more reliable estimate of long-term trajectory quality but incurs higher computation cost, whereas a smaller $L$ reduces overhead but may fail to capture meaningful differences among candidates.
In practice, we find that moderate values of $L$ (e.g., $L=5$) are sufficient to balance these concerns, since code generation typically exhibits strong local dependencies and short-horizon lookahead already provides useful signals. We analyze the effect of varying $L$ through experiments in Section 4.5.
4. Evaluation
@@
  • RQ3 (Ablation Study): How effective is the logistic regression–based threshold learning component, and can the learned thresholds better guide practical decoding?
  • RQ4 (Parameter Sensitivity): How does AdaDec perform under different values of the lookahead length $L$?
4.1. Experimental Setup
Our experimental setup is as follows.
@@ 4.1. Experimental Setup
Datasets. We evaluate our method on three widely used code generation benchmarks:
  • HumanEval+ (Liu et al., 2023): An extended version of the original HumanEval(Chen et al., 2021b) dataset, comprising 164 hand-crafted Python programming problems. The number of test cases has been increased by nearly 81×, enabling a more comprehensive evaluation of the functional correctness of generated code.
  • MBPP+ (Liu et al., 2023): A refined and optimized subset of the original MBPP(Austin et al., 2021) dataset, comprising 378 Python programming problems. MBPP+ reduces the number of problems compared to MBPP but increases the size of the unit tests by 35×, providing a much stricter evaluation of the functional correctness of LLM-generated code.
  • DevEval (Li et al., 2024): A repository-level benchmark aligned with real-world code repositories. It includes 1,874 testing samples from 117 repositories across 10 domains, providing a comprehensive evaluation of code generation models in practical software development scenarios. In our experiments, we randomly sample 238 problems. This sample size follows the statistical guideline that 238 measurements are sufficient to ensure, with 90% confidence, that the estimated performance lies within a ±5% margin of the true value.
Baselines. We compare AdaDec against two standard decoding strategies:
  • Greedy Decoding (Radford et al., 2019): A straightforward approach where at each step, the model selects the token with the highest probability as the next token. This method is computationally efficient but may lead to suboptimal solutions due to its myopic nature.
  • AdapT (Zhu et al., 2023): An uncertainty-guided adaptive decoding strategy that adjusts the sampling temperature for tokens at the beginning of a line (high-uncertainty tokens) to increase randomness and help overcome difficult-to-predict points, while keeping low-uncertainty tokens more deterministic. For our experiments, we set the parameters as follows: $a=0.5$ for high-uncertainty tokens and $b=0.05$ for all other tokens.
We also considered Uncert-CoT (Zhu et al., 2025), another uncertainty-guided decoding method originally proposed for chain-of-thought reasoning tasks. However, we do not include it in our experimental comparison since no official implementation is available and the original paper omits several crucial details necessary for faithful reproduction.
4.2. RQ1: Effectiveness in Improving Functional Correctness
@@
Table 2 presents the Pass@1 performance of three decoding strategies, Greedy, AdapT, and AdaDec, across seven LLMs evaluated on the HumanEval+, MBPP+ and DevEval datasets.
@@
Table 2. Pass@1 comparison of Greedy, AdapT, and AdaDec on HumanEval+, MBPP+, and DevEval. Values in parentheses indicate the change relative to Greedy.
Model HumanEval+ MBPP+ DevEval
Greedy AdapT AdaDec Greedy AdapT AdaDec Greedy AdapT AdaDec
DS-1.3B 51.83% 52.44%(+0.61%) 57.32%(+5.49%) 41.01% 41.27%(+0.26%) 49.47%(+8.46%) 13.45% 13.03%(-0.42%) 18.49%(+5.04%)
DS-6.7B 62.80% 63.41%(+0.61%) 65.24%(+2.44%) 58.47% 58.73%(+0.26%) 61.11%(+2.64%) 28.57% 28.57%(0.00%) 27.73%(-0.84%)
ST-3B 48.78% 51.83%(+3.05%) 50.61%(+1.83%) 38.89% 39.42%(+0.53%) 46.03%(+7.14%) 15.55% 13.03%(-2.52%) 22.69%(+7.14%)
QW-0.6B 17.07% 18.29%(+1.22%) 22.56%(+5.49%) 9.52% 11.11%(+1.59%) 30.42%(+20.90%) 2.94% 2.52%(-0.42%) 7.56%(+4.62%)
QW-1.7B 40.24% 42.07%(+1.83%) 44.51%(+4.27%) 37.57% 36.51%(-1.06%) 42.86%(+5.29%) 7.98% 7.56%(-0.42%) 11.76%(+3.78%)
QW-4B 53.05% 53.66%(+0.61%) 56.10%(+3.05%) 40.48% 41.01%(+0.53%) 51.32%(+10.84%) 9.24% 9.24%(0.0%) 16.81%(+7.57%)
QW-8B 56.71% 57.32%(+0.61%) 62.80%(+6.09%) 29.37% 29.63%(+0.26%) 47.09%(+17.72%) 4.62% 4.20%(-0.42%) 7.56%(+2.94%)
AVG. $\Delta$ +1.22% +4.09% +0.34% +10.43% -0.60% +4.32%
Comparison with Greedy Search. Across all three benchmarks, AdaDec consistently outperforms Greedy decoding across most datasets and models. On HumanEval+, AdaDec improves pass@1 by an average of +4.09%, with gains as high as +6.09% for QW-8B and +5.49% for DS-1.3B. On MBPP+, the improvement is even more pronounced: AdaDec yields an average gain of +10.43%, with the most substantial boost observed for QW-0.6B (+20.90%) and QW-8B (+17.72%). On DevEval, AdaDec also yields meaningful improvements for most models (+4.32% on average), with ST-3B and QW-4B showing particularly strong gains (+7.14% and +7.57%, respectively). The only exception is DS-6.7B, where we observe a slight drop ($0.84\%$). This may be due to DevEval’s repository-level nature, where future logic are harder to capture with short-horizon lookahead, and DS-6.7B already performs strongly under Greedy decoding, leaving less room for improvement.
Comparison with AdapT. When compared to AdapT, AdaDec again shows clear advantages. While AdapT offers marginal improvements over Greedy decoding (e.g., +1.22% on HumanEval+ and +0.34% on MBPP+), it sometimes underperforms (e.g., -0.60% on DevEval). In contrast, AdaDec consistently achieves larger gains, surpassing AdapT by wide margins across nearly all models and datasets. For example, on MBPP+, AdaDec exceeds AdapT by more than +10% on average, with the gap reaching +19.31% for QW-0.6B. On HumanEval+, AdaDec achieves substantially higher gains (average +4.09% vs. +1.22%), highlighting its stronger effectiveness in enhancing code correctness.
Summary: Overall, the results indicate that AdaDec is highly effective in improving the functional correctness of generated code. Compared with Greedy search, it achieves consistent and often substantial improvements across most models and benchmarks. Compared with AdapT, AdaDec demonstrates more robust and scalable gains. These findings confirm the effectiveness of the proposed approach in enhancing decoding quality.
4.3. RQ2: Efficiency in Computational Cost
@@
We evaluate the efficiency of AdaDec in terms of computational cost on the HumanEval+, MBPP+ and DevEval datasets from the following perspectives. First, we measure the pause rate of AdaDec to assess the possible overhead introduced by its pause-then-rank mechanism. Second, we compare the prediction latency of AdaDec against Greedy decoding and AdapT method.
@@
Pause Rate. As presented in Table 3, the pause rate ranges from 4.80% to 12.17%, with an average of 7.59%. This low activation frequency indicates that AdaDec concentrates computational effort on a small subset of high-uncertainty steps, effectively balancing local correction with overall efficiency. Across all datasets, the pause rate remains below 10% for all models except QW-1.7B, demonstrating that selective triggering successfully avoids the high cost associated with global lookahead.
Table 3. Pause ratio of the models on HumanEval+, MBPP+, and DevEval.
DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B
HumanEval+ 6.51% 5.48% 4.80% 6.11% 11.17% 7.82% 5.98%
MBPP+ 8.90% 7.46% 6.52% 8.50% 12.17% 7.11% 8.33%
DevEval 7.41% 6.17% 6.59% 6.16% 11.47% 9.69% 5.04%
AVG. $\Delta$ 7.61% 6.37% 5.97% 6.92% 11.60% 8.21% 6.45%
Latency. Table 4 reports the average generation time per problem for Greedy, AdapT, and AdaDec across HumanEval+, MBPP+, and DevEval. Overall, AdaDec introduces moderate additional latency compared to Greedy, reflecting the computational overhead of the pause-and-rerank mechanism. On most models and datasets, the increase remains reasonable—for instance, +0.23s on HumanEval+ and -0.84s on MBPP+ on average—demonstrating that selective triggering effectively balances enhanced generation quality with efficiency.
AdapT exhibits mixed runtime behavior. In some cases, it is faster than Greedy (e.g., DS-1.3B on MBPP+), likely because higher-temperature sampling applied only at high-uncertainty tokens (typically line-start tokens) can lead to shorter, simpler continuations, reducing the total number of decoding steps. This effect is particularly noticeable on smaller models and in datasets like DevEval. In contrast, AdaDec performs targeted lookahead at every high-uncertainty step, which increases runtime, especially for larger models and more complex DevEval problems, but provides more reliable improvements in functional correctness.